OpenStudy (anonymous):

what is 8 times the square root of 54 - 4 times the square root of 6

OpenStudy (saifoo.khan):

\[(8 \times \sqrt{54}) - ( 4 \times \sqrt6)\]This?

OpenStudy (anonymous):

\[8\sqrt{54}-4\sqrt{6}\] \[8\sqrt{9*6}-4\sqrt{6}\] \[24\sqrt{6}-4\sqrt{6}\] \[20\sqrt{6}\]

OpenStudy (anonymous):

not really, what kushashwa23 put

OpenStudy (saifoo.khan):

o.O

OpenStudy (anonymous):

have u got it

OpenStudy (anonymous):

yeah thanks :)